# Thinking beyond the anthropomorphic paradigm benefits LLM research

### v1

@@ 5 Recommendations
Develop new conceptualizations that capture the distinct nature of LLMs
Although our analysis highlights the limitations of anthropomorphic assumptions, it does not negate the value of drawing on human psychology and cognitive science, especially for developing new metaphors through which to make sense of LLMs as distinct systems. For example, McCoy et al. (2023)’s “teleological approach”, while drawn from cognitive science, is used to illuminate fundamental differences in how humans and LLMs operate and ought to be evaluated. Additionally, Shanahan et al. (2023) “role-play” conceptualization (and similarly Andreas (2022) “agent models”), while employing folk psychological terms, does so with conceptual precision that clarifies our understanding of LLMs as unique systems.
@@
There are several fields that offer valuable, less anthropocentric frameworks for current challenges in developing, aligning, and evaluating LLMs. For example, systems engineering and control theory provide established frameworks for analyzing feedback loops between LLMs and their environment (e.g., users) and understanding in-context alignment challenges (Pan et al., 2024). Similarly, design studies and HCI research offer theoretical (e.g., affordance theory) and methodological (e.g., user studies) tools for improving human interactions with increasingly social and human-like systems (Ibrahim et al., 2024). Software engineering offers proven methods for building and testing reliable systems, such as those with LLM-based agents, at scale (Bass et al., 2025).
@@ 6 Alternative views
Anthropomorphic conceptualizations of LLMs serve important purposes. They provide intuitive frameworks for understanding complex systems, offer pragmatic terminology for discussing model behavior, and have historically driven significant advances in the field. Many researchers argue that these human-centered analogies are not just convenient, but essential for developing and deploying AI systems that will ultimately interact with humans.
Critics of our position might also note that anthropomorphic thinking has led to breakthroughs in LLM research. They might argue that human cognition provides a proven template for intelligence, making it a valuable guide for AI development. And, that since LLMs are trained on human-generated data and designed to interact with humans, some degree of anthropomorphic framing may be inevitable and even desirable.
Throughout this position paper, we have acknowledged the alternative view that anthropomorphic conceptualizations are (1) natural and pragmatic, as well as (2) helpful. We do not advocate for eliminating anthropomorphic concepts entirely. Rather, we suggest that expanding beyond purely anthropomorphic frameworks can reveal new and potentially clarifying research directions. Our critique focuses specifically on how certain anthropomorphic assumptions may constrain research questions and methodologies. We also do not engage in debates about LLM sentience, agency, or reasoning capabilities. Instead, we examine how assumptions about human-like characteristics shape — and potentially limit — our understanding of these systems.
Impact Statement
This paper primarily addresses the computer science research community, examining research priorities and methodologies. We acknowledge that anthropomorphic terminologies and concepts also shape the broader public’s perceptions, usage, and likely regulation of these systems (Gilardi et al., 2024). As AI research gains broader visibility, its framing and communication may impact public understanding beyond this position paper’s scope. We encourage both future work and readers of this work to consider impacts on the public beyond those discussed here.

### v2

@@ 5 Recommendations
Develop new concepts & metaphors that capture the distinct nature of LLMs
Although our analysis highlights the limitations of anthropomorphic assumptions, it does not negate the value of drawing on human psychology and cognitive science, especially for developing new metaphors through which to make sense of LLMs as distinct systems. For example, McCoy et al. [59]’s “teleological approach", while drawn from cognitive science, is used to illuminate fundamental differences in how humans and LLMs operate and ought to be evaluated. Additionally, Shanahan et al. [79] “role-play" metaphor (and similarly Andreas [4] “agent models"), while employing folk psychological terms, does so with conceptual precision that clarifies our understanding of LLMs as unique systems.
@@
There are several fields that offer valuable, less anthropocentric frameworks for current challenges in developing, aligning, and evaluating LLMs. For example, systems engineering and control theory provide established frameworks for analyzing feedback loops between LLMs and their environment (e.g., users) and understanding in-context alignment challenges [73]. Similarly, design studies and HCI offer theoretical (e.g., affordance theory) and methodological (e.g., user studies) tools for improving human interactions with increasingly social and human-like systems [40]. Software engineering offers proven methods for building and testing reliable systems, such as those with LLM-based agents, at scale [11].
@@ 6 Alternative views
Anthropomorphism in LLM research serves important technical and sociotechnical purposes. On the technical side, it provides intuitive frameworks for understanding complex systems and offers pragmatic terminology for discussing model behavior. Some might argue that human cognition provides a proven template for intelligence, making it a valuable guide for AI development that has already led to breakthroughs. And, since LLMs are trained on human-generated data and designed to interact with humans, some degree of anthropomorphic framing may be inevitable and even desirable. On the sociotechnical side, anthropomorphic framing might improve our ability to engage non-technical stakeholders, communicate ethical considerations, and support policy discussions.
Throughout this position paper, we have acknowledged the alternative view that anthropomorphic thinking is (1) natural and pragmatic, as well as (2) helpful. We do not advocate for eliminating it entirely. Rather, we suggest that awareness of the prevalence and limitations of anthropomorphic thinking can reveal new and potentially clarifying research directions. Our critique focuses specifically on how certain anthropomorphic assumptions may constrain research questions and methodologies. Further, we acknowledge the role of anthropomorphic framing in communicating with the public [21]. While it is valuable, we also draw attention to the fact that it can result in public misconceptions and reduced AI literacy, for example, when systems are conceptualized as having human-like agency and power that they do not possess.